solve.
-5/3 x = - 3/5

Answer by ms.nelson(2)

This is not a quadratic equation.
When give a problem like this with a fraction attached to a variable on one side and a constant (a number with no variable) on the other you can do the problem in one step by multiplying by the reciprocal of the fraction that is attached to the variable.
Reciprocals of fractions are just the fractions flipped. For example the reciprocal of 1/2 is 2/1. The reciprocal of - 2/3 is -3/2. If it was negative starting out then it will be negative at the end and the same goes for positive.
-5/3x= -3/5
The reciprocal of -5/3 is -3/5....
multiply both sides by the reciprocal
-5/3x = -3/5
-3/5(-5/3)x = -3/5(-3/5)(multiply -3/5 by both sides)
x = 9/25
check your solution by substituting 9/25 for x in the original and see if it works
-5/3 (9/25) =? -3/5
-3/5 = -3/5 it works, there fore x = 9/25 is correct